OpenStudy (chris215):

What is the instantaneous slope of y = negative 7/x at x = 3?

OpenStudy (tkhunny):

Shall we utilize the 1st Derivative?

OpenStudy (anonymous):

tkbunny is right, you need to solve this by first finding the derivative of \[-\frac{ 7 }{ x }\] and then plug 3 in your new formula

imqwerty (imqwerty):

LOL *tkhunny

OpenStudy (anonymous):

eh bunny sounds hotter

imqwerty (imqwerty):

yeah xD

imqwerty (imqwerty):

slope is given by this expression -\[slope=\frac{ dy }{ dx }\]

OpenStudy (tkhunny):

It may be an algebra exercise. I'm still waiting for a hint from the OP.

imqwerty (imqwerty):

@chris215

OpenStudy (chris215):

is the derivative 7x^-2?

imqwerty (imqwerty):

u can write 7/x like this-\[7x^{-1}\] :)

OpenStudy (anonymous):

When possible, you want to get rid of the fraction and that means you want to move x to the numerator. And how is that done? Remember that \[x = x ^{1}\] because any number raised to the power of 1 is simply that number, and when using derivatives, you want to make all your x's in exponential formats Then you can rewrite it like this 

OpenStudy (anonymous):

yep

imqwerty (imqwerty):

\[\frac{ 1 }{ a^b }=a^{-b}\]

OpenStudy (anonymous):

then you can move the x^-2 to the bottom to make it a positive exponent, plug 3 in the "x" and then that will give you the instantaneous slope

OpenStudy (anonymous):

sorry it's actually -7x^-2

OpenStudy (anonymous):

because using the exponent rule for derivatives